OpenStudy (anonymous):

Describe how the graph of y = x2 can be transformed to the graph of the given equation. y = (x+4)2

OpenStudy (aum):

f(x) = x^2 f(x+4) = (x+4)^2 Changing f(x) to f(x+4) is equivalent to shifting the graph of f(x) to the LEFT by 4 units.

OpenStudy (anonymous):

thank you so much